OpenStudy (australopithecus):

Determine and show if this series is convergent or divergent an = 4^(n+3)/3^(n+4) so I have 4^(3) (lim 4^(n)) / 3^(4)(lim 3^(n)) n->infinity n->infinity How would I compute the limit?

OpenStudy (australopithecus):

it seems to me that it would be 4^(3)/3^(4) but wolfram alpha is giving me zero as the limit anyone able to explain to me why that is?

OpenStudy (australopithecus):

wait it displays monotonicity (4^(3)/3^(4)) lim (4/3)^(n) n-> infinity

OpenStudy (australopithecus):

I cant use squeeze law on it though

OpenStudy (australopithecus):

because there is no upper bound

OpenStudy (ash2326):

Find \(\Large \frac{a_{n+1}}{a_n}\), if it's a constant, then the series will converge. Could you check it?

OpenStudy (australopithecus):

(4^(n+1)/3^(n+1))/4^(n)/3^(n) = 4^(n+1)3^(n)/3^(n+1)(4^(n)) = 4^(n)(4)(3^(n))/3^(n)(3)(4^(n)) = 4/3

OpenStudy (australopithecus):

what rule is that?

OpenStudy (australopithecus):

I can see the function is increasing but I dont think it has a lower bound or upper bound

OpenStudy (amistre64):

its the ration test\[\]

OpenStudy (amistre64):

ratio ....

OpenStudy (australopithecus):

I also have to compute the limit

OpenStudy (amistre64):

and using a convergence method that i have in my head (could be wrong tho)\[\lim_{n\to\ inf}\frac{4^{n+3}}{3^{n+4}}=\lim_{n\to\ inf}\frac{4^{n}4^{3}}{3^{n}3^{4}}\] \[K\ \lim_{n\to\ inf}(\frac{4}{3})^n\]

OpenStudy (australopithecus):

right thats what I have so far but I don't know where to go from there

OpenStudy (australopithecus):

also for that test to work doesn't it need to have an upper or lower bound?

OpenStudy (australopithecus):

It is pretty clear that the limit goes to 0 but how do I show that

OpenStudy (amistre64):

Lhopital perhaps?

OpenStudy (amistre64):

or integral rule?

OpenStudy (australopithecus):

do I use the rule Kln(L) = nln(3/4)

OpenStudy (australopithecus):

sorry I mean Kln(L) = K(nln(3/4))

OpenStudy (australopithecus):

then convert it into a fraction and use L'H

OpenStudy (amistre64):

dunno, that looks foreign to me 4^n/3^n is already a fraction; and I dont really see it Lhoping out

OpenStudy (australopithecus):

true, ok I will try that, well it is a rule apparently I have it in my notes

OpenStudy (amistre64):

derivatives will always end up with (4/3)^n

OpenStudy (amistre64):

any exponential function with a base greater then 1 diverges i think

OpenStudy (australopithecus):

I think I need to use that rule otherwise I get 4^(n)log(4)/3^(n)log(3)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=limit+4%5E%28x%2B3%29%2F3%5E%28x%2B4%29 perhaps, but you might wanna chk this out to chk things by

OpenStudy (australopithecus):

and it just will get more complex from there

OpenStudy (australopithecus):

weird it told me the limit was 0 when I put it into wolfram alpha

OpenStudy (amistre64):

n to -inf = 0 n to +inf = inf n to 0 = ....what it said

OpenStudy (australopithecus):

so it does go to infinity

OpenStudy (australopithecus):

so it is divergent

OpenStudy (amistre64):

correct, it diverges.

OpenStudy (australopithecus):

what is the test I can apply?

OpenStudy (amistre64):

ratio test should do it

OpenStudy (australopithecus):

doesnt it need an upper or lower bound though?

OpenStudy (amistre64):

thats the one ash presented at the start

OpenStudy (amistre64):

no

OpenStudy (amistre64):

ratio test looks at how fast the slope is changeing at the ends

OpenStudy (australopithecus):

so if an+1/an > 1 then it is divergent?

OpenStudy (amistre64):

lets see :) 4^(x+3) * 3^(x+3) ----------------- = 4/3 4^(x+2) * 3^(x+4) i think theres a different "rule"; but so far its seems right

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcII/RatioTest.aspx this should be more definitive than I can recall

OpenStudy (amistre64):

c < 1 < d if 1, then try different test

OpenStudy (australopithecus):

thanks

OpenStudy (amistre64):

yw